ASPEARS:

Rewrite the function to make it easy to graph using transformations of its parent function. Describe the graph. y=radical 4x-20

luhivqqcherry:

what are the choices?

ASPEARS:

there are none

ASPEARS:

how would i rewrite

Illuminous:

1 attachment
ASPEARS:

why is there a minus 5?

Illuminous:

1 attachment
luhivqqcherry:

-5

luhivqqcherry:

y=4x-20-5 simply: 7=2 x-5 -5

Illuminous:

W

Illuminous:

Team work ;-;

luhivqqcherry:

yess

ASPEARS:

thanks

ASPEARS:

i have one more

luhivqqcherry:

alr

ASPEARS:

same thing just different numbers. radical 25x-75

luhivqqcherry:

25x -75 then you have to factor it out 25 from the expression 25(x-3) your answer then would be 25(x - 3)

ASPEARS:

it said it was wrong there shoulr be a radical somewhere

luhivqqcherry:

oh.

Illuminous:

Original equation x = √25x-75 Isolate -√25x-75 = -x Organize. √25x-75 = x

luhivqqcherry:

x = 3.4861

Illuminous:

Raise both sides to the second power (√25x-75)2 = (x)2 After squaring 25x-75 = x2

Illuminous:

Rearranged equation x2 - 25x + 75 = 0 This equation has two real roots: x1 = ( 25 + sqrt(325) ) / 2 = 21.5139 x2 = ( 25 - sqrt(325) ) / 2 = 3.4861

Illuminous:

Original equation, root isolated, after clean up √25x-75 = x Plug in 21.5139 for x √25•(21.5139) -75 = (21.5139) Simplify √462.847 = 21.514 Solution checks !! Solution is: x = 21.5139

Illuminous:

Original equation, root isolated, after tidy up √25x-75 = x Plug in 3.4861 for x √25•( 3.4861) -75 = ( 3.4861) Simplify √12.153 = 3.486 Solution checks !! Solution is: x = 3.4861

Illuminous:

Which, makes your final answer x = 3.4861

Illuminous:

Okay, you can close post now.